OpenStudy (anonymous):

Jose received $30 for his birthday and put it in his piggy bank. Each week, he puts 4 more dollars in his bank. The amount of money in dollars D in his bank is given by D = 4w + 30 where w is the number of weeks he has saved 4 dollars. His sister Marisol makes a table of reasonable values for w. Marisol selects 16 as a value for w. What is the corresponding table entry for D? 34 434 50 94

OpenStudy (anonymous):

D=4w+30 D=4(16)+30 D=64+30 D=94 The answer is $94.

OpenStudy (anonymous):

thx love

OpenStudy (anonymous):

your welcome